# The Relational Origins of Rules in Online Communities

### v1

@@ 5.3.1. Reinforcing Community Identity
The process of customizing the adopted rule set revealed another motivation behind why rules exist: reinforcing community identity. While rules that structure community identity fulfill a functional purpose by restricting contributions that don’t achieve community goals, they also serve a relational purpose by defining the boundary between what the community is and what the community is not—setting a community apart from others that overlap in membership or topic. This was especially important for subreddits and Fandom wikis, where communities with similar topics attract the same audiences who may arrive with misaligned expectations about core community goals. For example, R4 adopted a rule requiring that contributions “Must be relevant to cybersecurity professionals” because newcomers would arrive asking for cybersecurity troubleshooting. R4 explained the rationale for this rule: “we’re trying to create a community for technical professionals.” Rules like these helped to make explicit the community’s intended audience, reinforcing the identity of their community. R4 further reinforced r/cybersecurity’s identity by creating an offshoot subreddit specifically designed for aggregating questions about cybersecurity help and directed newcomers seeking help to post there, linking to it in r/cybersecurity’s sidebar.
@@
In another example, R2 manages two subreddits: both involved on Q&A style discussions, but they differ in the level of formality and seriousness. R2 explained that the rules in r/explainlikeimfive have to be strict and structure a “rigorous environment” while the rules in r/nostupidquestions allow for more freedom in what can be asked and how it can be answered, noting that the rules help keep that community’s discussions “as open as humanly possible.” R2 said “We [at r/nostupidquestions] want a place that’s not r/explainlikeimfive.” By establishing these distinct expectations with their rules, both communities thrived despite overlapping a similar niche space. The rules helped distinguish the communities by the style of Q&A format: one for formal and serious discussions and one that’s completely informal and casual. The tailored rules communicate to both members and newcomers not only what the community is, but also what it is not.
@@
Rules also helped reinforce identity in Fandom communities. For example, FA9, a leader for the DC Wiki (a fandom for the comic book franchise), compared their rules to the rules of the Marvel Wiki (a fandom for a different comic book franchise) to explain how their community was unique. Both wikis utilized a Fandom feature, which adds a forum for members to have discussions about their wiki’s content. That said, FA9 noted that in the DC Wiki, rules there promote “meaningful discussions” by restricting content related to fan fiction, while the Marvel Wiki has more “permissive” rules, allowing its members to discuss a variety of things, including fan fiction. FA9 lamented the disruption from newcomers expecting rules like those in the Marvel Wiki’s community, telling us that the DC Wiki is “not the place” for that kind of content. Like R2, FA9 used rules to explain what their community is and isn’t about: members are supposed to stick closely to discussing content published in the comic book franchise and nothing else, including other media sources and fan fiction.
@@
Identity construction from rules was less explicit coming from Fediverse leaders. While some communities portrayed an identity (e.g., a Mastodon server for people in the Bay Area of California or a community for amateur radio enthusiasts), their rules usually did not limit discussions based on their community’s identity. However, some Fediverse leaders repeatedly mentioned that rules restricting discrimination and hate speech helped to signal a virtue of intolerance to intolerance. In other words, some Fediverse leaders saw the absence of such rules in other communities as an indicator that they allow hate speech to flourish. There were called “free speech” servers because it was inferred that content moderation was virtually absent. FV13 said leaders of free speech communities “don’t care about moderation–[they let you] say what you like.” FV8 told us that they adopted their strict rules on discrimination and hate speech to “signpost” to outsiders that their community is not a place for “free speech warriors.” In one instance, FV14 said their rules essentially help people self-select out because registration requires them to read their anti-discrimination rules.
@@ 5.4. Encountering Problems that Challenge the Rules
When asked why the rules were changed, leaders tended to give functional accounts by describing how various problems would raise questions of whether adopted rules sufficed to achieve community goals. However, not all communities experienced such problems or felt the need to change rules at all—particularly those closed to new members or low in activity. For FV17, an absence of problems in their community resulted from the tight-knit and homophilous nature of the group who “kind of know each other already”. Notably, on the Fediverse, leaders can close account registrations on their server and thus control who may become a member. For example, FV10’s server “pretty much never had open registration” and thus never had to change rules. Similarly, relatively small or new communities, like FA6’s wiki community, rarely saw problems with rule breaking because “the number of edits was in the very low volume.” When membership and activity are low, there are less opportunities for problems to emerge and threaten community goals. But even if a community has not experienced problems, this does not mean it never will. Our interviewees discussed how a community might encounter problems—and then change its rules in response to them—in varying ways over the community’s lifecycle, particularly as communities grew or declined.
@@ 5.4.1. Problems from Growth
As communities grew in membership and activity, leaders increasingly had to consider whether the community is still able to achieve its goals under the old rule set. In particular, as communities grow and the number of people attempting to contribute increases, content-related issues tend to emerge and occasion leaders to reevaluate and reinforce their communities’ goals by making rules controlling quality and identity.
@@
Reliability is an important feature of many wikis because people use them for acquiring information about a specific topic. Allowing the creation of articles that aren’t grounded in factual basis might jeopardize the goal of being a reliable source of information.
Community growth also tended to coincide with the addition of rules restricting spam. FV5 explained how in their Fediverse server for musicians in Boston, they added a rule to control spam by self-promoting local artists:
@@
Even though the problem was caused by good-faith local artists promoting their shows, FV5 felt that too many such posts threatened the community’s goal “to be a space for people to talk each other.” Similarly, R5’s community, a news discussion space for a region in Alabama, created a new rule to limit posts about topics that gain sudden virality and importance because they can set off a cascade of members posting threads about the same news article or topic. R5 explained their rationale for rulemaking to prevent such spam during the COVID-19 pandemic:
@@
Implementing a rule restricting posts about suddenly important topics to a single “Mega Thread” helped to reduce the noise that was generated by the sudden influx of people posting the same news articles and allow space for other topics related to their community’s goals.
@@ 5.4.2. Problems from Decline
Leaders also attributed rule change to community decline manifesting in diminishing member activity or neglecting members’ problems. Some leaders responded to periods of diminishing member activity by changing the rules to reduce barriers to participation. In FA5’s community, members had complained on Talk Pages about “feeling kind of suffocated” by the rules’ requirements for contributing. The leaders listened and made these requirements less restrictive believing that they “need to not be so hardline on handling these things.” FV5 told us how they hadn’t changed rules yet but “might have to take rules away if they seem overly restrictive” for members of their Fediverse server. A more extreme example of changing the rules due to diminishing member activity comes from R3’s community, a place to discuss a popular US science-fiction TV series. Following the show’s conclusion, activity in the subreddit began to plummet. According to R3, their community’s moderators responded problem by evaluating not just the rules, but the community’s identity:
@@
“Relaxing the rules” to allow fan-fiction and memes would open up new opportunities for members to make novel contributions to the community and sustain activity that isn’t solely rehashing discussions about the show’s content. That said, R3’s leaders also brought in to question what kind of community they would be: Would they preserve their exclusive focus on discussing the show’s content or become a more casual space with fan stories and low-effort memes?
@@
In rarer cases, communities recovered from neglecting members’ problems in the rules only by radically changing leadership first. For example, FA12 told us that the previous leadership team of their Fandom wiki was changed for being “lax with users who were being discriminatory, which was very disheartening to see for a while.” Harassment on wiki Talk Pages went unaddressed by moderators, leaving members feeling unsafe. The problem was eventually addressed through a “big overhaul of the moderator team” who also updated their rules to crack down on discriminatory language. In R9’s community, where members exchange plants through the postal mail, the previous leadership was “pushed out because they weren’t putting in any rules to protect people from scams.” As in FA12’s example, members did not feel safe and participation declined as the leaders failed to protect them. R9, along with some other core members, became their community’s new leaders through a community-wide vote and immediately added rules requiring verification photos, commenting on posts before privately messaging, and using PayPal Goods and Services for plant trades. Both of these cases show that communities may feel the need to completely change their leadership in order to survive periods of decline, especially if leaders overseeing the problems causing decline cannot or will not change the rules themselves.
@@ 5.5. Becoming Aware of the Need to Change the Rules
While problems themselves are primarily functional matters, we found that each stage in the processes of changing rules has significant relational aspects. The first stage of a rule change, however minor or major, is becoming aware of (new) problems so as to recognize the need to change, add, or delete as a rule. Here we elaborate two ways community leaders became aware of problems.
@@ 5.5.1. Reporting to Leadership
Community leaders often learned about problems not through direct experience, but from fellow leaders or community members. Leaders in FA1’s Fandom wiki, having recognized a problem, would “always discuss it with the [other] moderators,” to decide on whether they should implement a change. In many communities the leadership relied on members for awareness of problems. Rule change processes would happen “whenever the user base complains” (FA11), in which“someone has an issue with a rule, they bring it up—on the forum—to an admin” (FA3). Subreddits’ leaders learned about problems from members not only through direct messages or the ModMail system, but also in “meta” threads for conversations about the subreddit. Some subreddit leaders openly encouraged such feedback to learn how to improve governance. Still, some problems were acknowledged by leaders but left unaddressed. Leaders explained that their limited capacity as volunteers constrained their ability to find solutions to problems. For example, R1’s community leaders “never really had the time” to improve governance in their community because they were “kind of busy with real life.” It wasn’t until the COVID-19 pandemic that R1’s community leaders, either laid off from work or working from home, found extra time to spend on improving their community’s rules.
@@ 5.5.2. Reflecting After Enforcement
Our interviews highlighted how everyday practices of enforcing governance in the community could prompt reflection about rules after enforcement actions. For example, FA3 said that their community lacked rules in the early days and was like “the wild west.” The problems that arose for them eventually led them to craft rules that would address the behavior and content issues they had noticed. While FA3 points to an example where new rules were motivated by absence of rules, R8 noted how the enforcement of existing rules could reveal ways the current rules were insufficient. This would ultimately lead to discussions where community leaders took stock of their rule sets and reflected on how rules could be improved:
@@
Our interviews suggest that these moments of reflection often involved formalizing and streamlining norms that emerged. FA4 explained that a reason for one of their biggest rule changes was “to overhaul” their communities’ rules to “properly consolidate all the unwritten rules” that they had accumulated. Leaders may also overhaul rule sets when they become too bloated with redundant or unnecessary rules that were also hard to enforce. FA12 explained they overhauled rules this way because their rule set was “really complicated and kind of difficult to keep track of.” In sum, sweeping rule changes usually occurred in mature communities where inaction would threaten community goals.
@@ 5.6. Deliberating on How to Solve Rule-Related Problems
When implementing a solution isn’t obvious or may disrupt community members’ experiences, leaders often engage in a deliberation process to better make sense of their current problem, what the current rules do to combat the problem, and how possible changes might ultimately resolve the problem. Interviews revealed how moderators and admins would create meeting spaces to discuss problems and how to handle them. Leaders on Fandom and Reddit often convened in group chats using software such as Discord where “everybody [is] adding their own [two] cents to it and then changing [rules] as a group.”(R7) FA1 described a situation where their team of moderators got on a voice call for over 3 hours reviewing and “cleaning” up their rule pages. FA4 similarly recalled a “lengthy process with the rest of the administrative team” when they drafted new rules for their Fandom wiki. Leaders described various procedures for these discussions. For example, R9 said that they want to make sure the other leaders “all get along well” so that they can “communicate disagreements and not have it be a big thing.” When disagreements did occur, the leaders were “respectful enough that they’ll get to a point where one of them is swayed one way or the other, or [the issue] is just tabled for another month or two.” Other procedures for rule change discussions involved teasing out potential issues. FA6 described this process:
@@
“We try to get input from other users, and if we come to a decision and everyone else is like ’No, that decision sucks’, then we’re like ’Well we think it works better because this, this, and this, but we see your point and we can compromise here, here, and here on that change.”
In this example, FA3’s community worked collaboratively on changing rules by finding compromises between the goals of the leaders and the experiences of community members. R4 explained that this approach to rule change involved “consultation with the community.” For example, community members would make problems visible to leaders by “flagging posts” and “complaining about this type of behavior.” Then, before implementing solutions, leaders would first communicate with members about their intended rule changes to ensure the changes wouldn’t disrupt their experiences in the community. Leaders on Reddit in particular would take into account the “upvote” feature that allows community members to vote on particular comments within meta discussion threads to help leaders understand “what people were wanting” (R5).
@@ 5.7. Making the Decision to Change the Rules
Finally, after encountering and observing problems and deliberating over potential solutions, community leaders needed to ultimately decide whether or not to change the rules. We consider this a separate step from deliberation because communities frequently differed in their decision-making procedures, which reflected differences in governance philosophies and division of labor. Governance procedures vary to a degree within each setting, but we observed some patterns: 1) autocratic decision-making for leaders of Fediverse communities, 2) oligarchic decision-making for leaders on Fandom, and 3) a (limited) implementation of democratic participation for rule change decision-making for some Reddit communities.
@@
Decision-making processes were more involved for communities on Fandom and Reddit. In both platforms, communities were typically governed by more than one person. For Fandom communities, decision-making was typically “oligarchical”—decisions for a wiki community would be made within “a small collective of moderators and admins” (FA3). Leaders often sought consensus on how to handle problems and in discussions on changing the rules: “You always try to compromise or try to make your interests—make your points—to have these rules [or] not have these rules” (FA14). That said, we see hierarchical decision-making structures even within these small groups of leaders. FA1 pointed out that the original founder of their wiki community changed their rules with no deliberation from others when they came back from a hiatus “for half a year” because they “decided that the rules were too lax.” Fandom wikis’ hierarchies were often formalized in roles including bureaucrats (who held the most decision-making power), admins (who helped manage community-wide problems), and moderators (who did much of the rule enforcement). In FA11’s wiki, the “main bureaucrat” is “the final decision maker” who has “the central role of helping make the rules for the wiki and making the final decisions if there was a tie” when voting on rule changes.
@@
“We suggested a couple options that we’re like ‘Just remove them all’ or ‘Keep them but flair [^2] Flairing a post means that the title of the discussion post is labeled with a tag, in this case it would be “Advertisement.” them and mark them as advertising’, that kind of stuff. And that ended up being the result that we decided was probably best for the community. That’s forever another rule.”
@@
R9’s community had a history of neglect by former leaders who failed to implement rules that protect their members. When R9 and other leaders replaced the old guard, they decided to take an approach that incorporated community members into the decision-making process for changes. R9 explained that they do “community posts” once a month to allow members to give feedback about their experiences. When reviewing the feedback, if R9 sees “people a little bit cranky about some rules” that they have, they will “put out a poll…to see if they [members] want things to change or not.” Members vote on whether a change should be implemented, although sometimes, R9, told us, “many polls…end up basically being 50/50.” They shelve the problem and do another poll a month later. In this way, R9’s leaders ensure that they “change things as the community wants to be changed,” in order to make sure the community is going in the right direction, supporting community members’ goals, and safeguarding their experiences.

### v2

@@ 5.3.1. Reinforcing Community Identity
The process of customizing the adopted rule set revealed another reason that rules exist: reinforcing community identity. While rules that structure community identity fulfill a functional purpose by restricting contributions that don’t achieve community goals, they also serve a relational purpose by setting a community apart from others that overlap in membership or topic. This was especially important for subreddits and Fandom wikis, where overlapping communities attract audiences who may arrive with misaligned expectations from their experience in similar groups. For example, R4 adopted a rule requiring that contributions “must be relevant” to IT professionals because newcomers would arrive asking for IT troubleshooting. R4 explained the rationale for this rule: “we’re trying to create a community for technical professionals.” Rules like these helped to make explicit the community’s intended audience, reinforcing the identity of their community. R4 further reinforced their community’s identity by creating an offshoot subreddit specifically designed for aggregating questions about IT help and directed newcomers seeking help to post there.
@@
In another example, R2 manages two subreddits: both involved on Q&A style discussions, but they differ in the level of formality and seriousness. R2 explained that the rules in the first community have to be strict and structure a “rigorous environment” while the rules in the second community allow for more freedom in what can be asked and how it can be answered, noting that the rules help keep that community’s discussions “as open as humanly possible.” R2 said “We [at second community] want a place that’s not [first community].” By establishing these distinct expectations with their rules, both communities thrived despite overlapping a similar niche space. The rules helped distinguish the communities by the style of Q&A format: one for formal and serious discussions and one that’s completely informal and casual. The tailored rules communicate to both members and newcomers not only what the community is, but also what it is not.
@@
Rules also helped reinforce identity in Fandom communities. For example, FA9, a leader for another comic book franchise, compared their rules to the rules of the wiki for a different comic book franchise to explain how their community was unique. Both wikis utilized a Fandom forum feature. That said, FA9 noted that their wiki’s rules promote “meaningful discussions” by banning conversations about fan fiction, while the wiki they were contrasting theirs to has more “permissive” rules. FA9 lamented the disruption from newcomers expecting rules like those in the different wiki community, telling us that their wiki is “not the place” for that kind of content. Like R2, FA9 used rules to explain what their community is and isn’t about.
@@
Identity construction from rules was less explicit coming from Fediverse leaders. While some communities portrayed an identity (e.g., a Mastodon server for people in a specific city in the US or a community for amateurs of a specific hobby), their rules usually did not limit discussions based on their community’s identity. However, some Fediverse leaders repeatedly mentioned that rules restricting discrimination and hate speech helped to signal a virtue of intolerance to intolerance. In other words, some Fediverse leaders saw the absence of such rules in other communities as an indicator that they allow hate speech to flourish. These were called “free speech” servers and it was inferred that content moderation was virtually absent. FV13 said leaders of free speech communities “don’t care about moderation–[they let you] say what you like.” FV8 told us that they adopted their strict rules on discrimination and hate speech to “signpost” to outsiders that their community is not a place for “free speech warriors.” In one instance, FV14 said their rules essentially help people self-select out because registration requires them to read their anti-discrimination rules.
@@ 5.4. Encountering Problems that Challenge the Rules
Not all communities felt the need to change rules—particularly those closed to new members or low in activity. For FV17, an absence of problems in their community resulted from the tight-knit and homophilous nature of the group who “kind of know each other already”. Notably, on the Fediverse, leaders can close account registrations on their server and thus control who may become a member. For example, FV10’s server “pretty much never had open registration” and thus never had to change rules. Similarly, relatively small or new communities, like FA6’s wiki community, rarely saw problems with rule breaking because “the number of edits was in the very low volume.” When membership and activity are low, there are less opportunities for problems to emerge and threaten community goals.
But even if a community has not experienced problems, this does not mean it never will. Indeed, many communities had changed their rules in response to problems or challenges. Our interviewees discussed how a community might encounter problems—and then change its rules in response to them—in varying ways over the community’s lifecycle, particularly as communities grew or declined.
@@ 5.4.1. Problems from Growth
As communities grew in membership and activity, leaders increasingly had to consider whether the community was able to achieve its goals under the old rule set. In particular, as communities grow, content-related issues tend to emerge and occasion leaders to reevaluate and reinforce their communities’ goals by making rules controlling quality and identity.
@@
Reliability is an important feature of many wikis. Allowing the creation of articles that aren’t grounded in fact might jeopardize the goal of being a reliable source of information.
Community growth also tended to coincide with the addition of rules restricting spam. FV5 explained how in their Fediverse server for musicians in a large city, they added a rule to control spam by self-promoting local artists:
@@
Even though the problem was caused by good-faith local artists promoting their shows, FV5 felt that too many such posts threatened the community’s goal “to be a space for people to talk each other.” Similarly, R5’s community, a news discussion space for a city in the US , created a new rule to limit posts about topics that gain sudden virality and importance because they can set off a cascade of members posting threads about the same news article or topic. R5 explained their rationale for rulemaking to prevent such spam during the COVID-19 pandemic:
@@
Implementing a rule restricting posts about suddenly important topics to a single “Mega Thread” helped to reduce the noise of a sudden influx of people posting the same news articles and hold space for other topics related to their community’s goals.
@@ 5.4.2. Problems from Decline
Leaders also attributed rule change to community decline manifesting in diminishing member activity or neglecting members’ problems. Some leaders responded to periods of diminishing member activity by changing the rules to reduce barriers to participation. In FA5’s community, members had complained on Talk Pages about “feeling kind of suffocated” by rules. The leaders listened and made requirements less restrictive believing that they “need to not be so hardline on handling these things.” FV5 told us how they hadn’t changed rules yet but “might have to take rules away if they seem overly restrictive” for members of their Fediverse server. A more extreme example of changing the rules due to diminishing member activity comes from R3’s community, a place to discuss a television series. Following the show’s conclusion, activity in the subreddit began to plummet. According to R3, their community’s moderators responded by evaluating not just the rules, but the community’s identity:
@@
“Relaxing the rules” to allow fan-fiction and memes would open up new opportunities for members to make novel contributions to the community and sustain activity that isn’t solely rehashing discussions about the show’s content. That said, R3’s leaders also asked what kind of community they would be: Would they preserve their exclusive focus on discussing the show’s content or become a more casual space with fan stories and low-effort memes?
@@
In rarer cases, communities recovered from neglecting members’ problems in the rules only by radically changing leadership first. For example, FA12 told us that the previous leadership team of their Fandom wiki was changed for being “lax with users who were being discriminatory, which was very disheartening to see.” Harassment on Talk Pages went unaddressed by moderators, leaving members feeling unsafe. The problem was eventually addressed through a “big overhaul of the moderator team” who also updated their rules to crack down on discriminatory language. In R9’s community, where members exchange goods through the postal mail, the previous leadership was “pushed out because they weren’t putting in any rules to protect people from scams.” As in FA12’s example, members did not feel safe and participation declined as the leaders failed to protect them. R9, along with some other core members, became their community’s new leaders through a community-wide vote and immediately added rules requiring verification photos, commenting on posts before privately messaging, and using PayPal Goods and Services for plant trades.
@@ 5.5. Becoming Aware of the Need to Change the Rules
While rules addressing problems are primarily functional, we found that each stage in the processes of changing rules has significant relational aspects. The first stage of a rule change is always becoming aware of (new) problems so as to recognize the need to change, add, or delete as a rule. Here we elaborate two ways community leaders became aware of problems.
@@ 5.5.1. Reporting to Leadership
Community leaders often learned about problems not through direct experience, but from fellow leaders or community members. Leaders in FA1’s Fandom wiki, having recognized a problem, would “always discuss it with the [other] moderators,” to decide on whether they should implement a change. In many communities the leadership relied on members for awareness of problems. Rule change processes would happen “whenever the user base complains” (FA11), in which“someone has an issue with a rule, they bring it up—on the forum—to an admin” (FA3). Subreddits’ leaders learned about problems from members not only through direct messages or the ModMail system, but also in “meta” threads for conversations about the subreddit. Some subreddit leaders openly encouraged such feedback to learn how to improve governance. Still, some problems were acknowledged by leaders but left unaddressed.
@@ 5.5.2. Reflecting After Enforcement
Our interviews highlighted how everyday practices of enforcing governance in the community could prompt reflection about rules after enforcement actions. For example, FA3 said that their community lacked rules in the early days and was like “the wild west.” The problems that arose for them eventually led them to craft rules that would address the behavior and content issues they had noticed. While FA3 points to an example where new rules were motivated by the absence of rules, R8 noted how the enforcement of existing rules could reveal ways the current rules were insufficient. This would ultimately lead to discussions where community leaders took stock of their rule sets and reflected on how rules could be improved:
@@
Our interviews suggest that these moments of reflection often involved formalizing and streamlining norms. FA4 explained that a reason for one of their biggest rule changes was “to overhaul” their communities’ rules to “properly consolidate all the unwritten rules” that they had accumulated. Leaders may also overhaul rule sets when they become too bloated with redundant or unnecessary rules that were also hard to enforce. FA12 explained they overhauled rules this way because their rule set was “really complicated and kind of difficult to keep track of.” In sum, sweeping rule changes usually occurred in mature communities where inaction would threaten community goals.
@@ 5.6. Deliberating on How to Solve Rule-Related Problems
When implementing a solution isn’t obvious or may disrupt community members’ experiences, leaders often engage in a deliberation process to better make sense of their current problem, what the current rules do to combat the problem, and how possible changes might ultimately resolve the problem. Interviews revealed how moderators and admins would create meeting spaces to discuss problems and how to handle them. Leaders on Fandom and Reddit often convened in group chats using software such as Discord where “everybody [is] adding their own [two] cents to it and then changing [rules] as a group” (R7). FA1 described a situation where their team of moderators got on a voice call for over 3 hours reviewing and “cleaning” up their rule pages. FA4 similarly recalled a “lengthy process with the rest of the administrative team” when they drafted new rules for their Fandom wiki. Leaders described various procedures for these discussions. For example, R9 said that they want to make sure the other leaders “all get along well” so that they can “communicate disagreements and not have it be a big thing.” When disagreements did occur, the leaders were “respectful enough that they’ll get to a point where one of them is swayed one way or the other, or [the issue] is just tabled for another month or two.” Other procedures for rule change discussions involved teasing out potential issues. FA6 described this process:
@@
“We try to get input from other users, and if we come to a decision and everyone else is like ’No, that decision sucks’, then we’re like ‘Well we think it works better because this, this, and this, but we see your point and we can compromise here, here, and here on that change.”’
In this example, FA3’s community worked collaboratively on changing rules by finding compromises between the goals of the leaders and the experiences of community members. R4 explained that this approach to rule change involved “consultation with the community.” For example, community members would make problems visible to leaders by “flagging posts” and “complaining about this type of behavior.” Then, before implementing solutions, leaders would communicate with members about their intended rule changes to ensure the changes wouldn’t disrupt their experiences. Leaders on Reddit would take into account the “upvote” feature that allows community members to vote on particular comments within meta discussion threads to help leaders understand “what people were wanting” (R5).
@@ 5.7. Making the Decision to Change the Rules
Finally, after encountering and observing problems and deliberating over potential solutions, community leaders needed to ultimately decide whether or not to change the rules. Governance procedures vary to a degree within each setting, but we observed some patterns: 1) autocratic decision-making for leaders of Fediverse communities, 2) oligarchic decision-making for leaders on Fandom, and 3) a (limited) implementation of democratic participation for rule change decision-making for some Reddit communities.
@@
Decision-making processes were more involved for communities on Fandom and Reddit. In both platforms, communities were typically governed by more than one person. For Fandom communities, decision-making was typically “oligarchical”—decisions for a wiki community would be made within “a small collective of moderators and admins” (FA3). Leaders often sought consensus on how to handle problems and in discussions on changing the rules: “You always try to compromise or try to make your interests—make your points—to have these rules [or] not have these rules” (FA14). That said, we see hierarchical decision-making structures even within these small groups of leaders. FA1 pointed out that the original founder of their wiki community changed their rules with no deliberation from others when they came back from a hiatus “for half a year” because they “decided that the rules were too lax.” Fandom wikis’ leaders had formalized roles including bureaucrats (who held the most decision-making power), admins (who helped manage community-wide problems), and moderators (who did much of the rule enforcement). In FA11’s wiki, the “main bureaucrat” is “the final decision maker” who has “the central role of helping make the rules for the wiki and making the final decisions if there was a tie” when voting on rule changes.
@@
“We suggested a couple options that we’re like ‘Just remove them all’ or ‘Keep them but flair [^2] Flairing a post means that the title of the discussion post is labeled with a visible tag. them and mark them as advertising’, that kind of stuff. And that ended up being the result that we decided was probably best for the community. That’s forever another rule.”
@@
R9’s community had a history of neglect by former leaders. When R9 and other leaders replaced the old guard, they decided to take an approach that incorporated community members into the decision-making process for changes. R9 explained that they do “community posts” once a month to allow members to give feedback about their experiences. When reviewing the feedback, if R9 sees “people a little bit cranky about some rules” that they have, they will “put out a poll…to see if they [members] want things to change or not.” Members vote on whether a change should be implemented, although sometimes, R9, told us, “many polls…end up basically being 50/50.” They shelve the problem and do another poll a month later. In this way, R9’s leaders ensure that they “change things as the community wants to be changed,” in order to make sure the community is going in the right direction, supporting community members’ goals, and safeguarding their experiences.